# Can a MISL Fly? Analysis and Ingredients for Mutual Information Skill Learning

### v2

@@ B.4 Zero-shot goal reaching
Please see Fig. 6 for the full set of goal reaching results. We find CSF to generally perform closely to METRA, though slightly underperforming in Quadruped, Humanoid, and Kitchen. In Ant however, CSF outperforms METRA.
@@
We closely follow the setup in Park et al. (2024). For all baselines, 50 goals are randomly sampled from $[-50,50]$ in Ant, $[100,100]$ in HalfCheetah, $[-15,15]$ in Quadruped, and $[-10,10]$ in Humanoid. In Kitchen, we sample 50 times at random from the following built-in tasks: BottomBurner, LightSwitch, SlideCabinet, HingeCabinet, Microwave, and Kettle. In Robobin, we sample 50 times at random from the following built-in tasks: ReachLeft, ReachRight, PushFront, and PushBack.
@@ C.1 Quadratic approximation of $\textmd{LB}^{\beta}_{2}(\phi)$
We conduct experiments to study the accuracy of the quadratic approximation in Prop. 2 in practice. To answer this question, we reuse the METRA algorithm trained on the didactic Ant environment and compare $\log\mathbb{E}_{p(z)}[e^{(\phi(s^{\prime})-\phi(s))^{\top}z}]$ against $\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2}$. We can compute $\log\mathbb{E}_{p(z)}[e^{(\phi(s^{\prime})-\phi(s))^{\top}z}]$ analytically because $d=2$ in our experiments. Results in Fig. 9 shows a clear linear relationship between $\log\mathbb{E}_{p(z)}[e^{(\phi(s^{\prime})-\phi(s))^{\top}z}]$ and $\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2}$, suggesting that the slope of the least squares linear regression is near the theoretical prediction, i.e., $\lambda_{0}(d)=\frac{1}{2d}=0.25\approx 0.2309$. We conjecture that this linear relationship still exists for higher dimensional $d$ and, therefore, the second-order Taylor approximation proposed by Prop. 2 is practical.
@@ C.2 METRA and CSF are Sensitive to the Skill Dimension
METRA leverages different skill dimensions for different environments. This caused us to investigate what the impact of the skill dimension on exploration performance is. In Fig. 9, we find that both METRA (to a lesser extent) and CSF are quite sensitive to the skill dimension. We conclude that skill dimension is a key parameter to tune for practitioners when training their MISL algorithm.
@@ C.3 Sensitivity of CSF to the scaling coefficient $\xi$
Figure 10: The effect of $\xi$ in CSF. Increasing $\xi$ to slightly higher values (2, 5, 10) boosts performance of CSF, while a $\xi<1$ hurts performance substantially.
We conduct ablation experiments to study the effect of the scaling coefficient $\xi$ on the negative term of the contrastive lower bound ($\text{LB}_{-}^{\beta}(\phi)$ as well as investigating our theoretical prediction for selecting $\xi\geq 1$ in Appendix A.6. We compare the state coverage of different variants of CSF with $\xi$ choosing from $\{0.5,1,2,5,10\}$, plotting the mean and standard deviation over 5 random seeds. Results in Fig. 10 suggest that increasing $\xi$ to a higher values (2, 5, 10) boosts the state coverage of CSF, while a $\xi\leq 1$ hurts the performance. We choose to use $\xi=5$ in our benchmark experiments.
C.4 Sensitivity of METRA to $\lambda$, $\epsilon$, and the dual learning rate
@@ C.6 Failed Experiments
We experimented with both CSF and METRA on the MiniHack-Corridor-R5-v0 from MiniHack (Samvelyan et al., 2021). This environment has a discrete action space and requires the agent to navigate between 5 different rooms using various doors and corridors. None of these methods got consistently stable performance on this environment. We conjecture that the difficulties could come from (1) the discrete action space, (2) randomized initial states and dynamics (layout of the rooms), and (3) partial observability.

### v3

@@ B.4 Zero-shot goal reaching
Please see Fig. 6 for the full set of goal-reaching results. We find CSF to generally perform closely to METRA, though slightly underperforming in Quadruped, Humanoid, and Kitchen. In Ant, however, CSF outperforms METRA.
@@
We closely follow the setup in Park et al. (2024). For all baselines, 50 goals are randomly sampled from [50,50][-50,50] in Ant, $[100,100]$ in HalfCheetah, [15,15][-15,15] in Quadruped, and [10,10][-10,10] in Humanoid. In Kitchen, we sample 50 times at random from the following built-in tasks: BottomBurner, LightSwitch, SlideCabinet, HingeCabinet, Microwave, and Kettle. In Robobin, we sample 50 times at random from the following built-in tasks: ReachLeft, ReachRight, PushFront, and PushBack.
@@ C.1 Quadratic approximation of $\textmd{LB}^{\beta}_{2}(\phi)$
We conduct experiments to study the accuracy of the quadratic approximation in Prop. 2 in practice. To answer this question, we reuse the METRA algorithm trained on the didactic Ant environment and compare $\log\mathbb{E}_{p(z)}[e^{(\phi(s^{\prime})-\phi(s))^{\top}z}]$ against $\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2}$. We can compute $\log\mathbb{E}_{p(z)}[e^{(\phi(s^{\prime})-\phi(s))^{\top}z}]$ analytically because $d=2$ in our experiments. Results in Fig. 9 shows a clear linear relationship between $\log\mathbb{E}_{p(z)}[e^{(\phi(s^{\prime})-\phi(s))^{\top}z}]$ and $\lVert\phi(s^{\prime})-\phi(s)\rVert_{2}^{2}$, suggesting that the slope of the least squares linear regression is near the theoretical prediction, i.e., $\lambda_{0}(d)=\frac{1}{2d}=0.25\approx 0.2309$. We conjecture that this linear relationship still exists for higher-dimensional $d$ and, therefore, the second-order Taylor approximation proposed by Prop. 2 is practical.
@@ C.2 METRA and CSF are Sensitive to the Skill Dimension
METRA leverages different skill dimensions for different environments. This caused us to investigate the impact of the skill dimension on exploration performance. In Fig. 9, we find that both METRA (to a lesser extent) and CSF are quite sensitive to the skill dimension. We conclude that the skill dimension is a key parameter to tune for practitioners when training their MISL algorithm.
@@ C.3 Sensitivity of CSF to the scaling coefficient $\xi$
Figure 10: The effect of $\xi$ in CSF. Increasing $\xi$ to slightly higher values (2, 5, 10) boosts the performance of CSF, while a $\xi<1$ hurts performance substantially.
We conduct ablation experiments to study the effect of the scaling coefficient $\xi$ on the negative term of the contrastive lower bound ($\text{LB}_{-}^{\beta}(\phi)$ as well as investigating our theoretical prediction for selecting $\xi\geq 1$ in Appendix A.6. We compare the state coverage of different variants of CSF with $\xi$ choosing from $\{0.5,1,2,5,10\}$, plotting the mean and standard deviation over 5 random seeds. Results in Fig. 10 suggest that increasing $\xi$ to higher values (2, 5, 10) boosts the state coverage of CSF, while a $\xi\leq 1$ hurts the performance. We choose to use $\xi=5$ in our benchmark experiments.
C.4 Sensitivity of METRA to the dual variable, the slack, and the dual learning rate
@@ C.6 Failed Experiments
We experimented with both CSF and METRA on the MiniHack-Corridor-R5-v0 from MiniHack (Samvelyan et al., 2021). This environment has a discrete action space and requires the agent to navigate between 5 different rooms using various doors and corridors. None of these methods got consistently stable performance in this environment. We conjecture that the difficulties could come from (1) the discrete action space, (2) randomized initial states and dynamics (layout of the rooms), and (3) partial observability.